# OmniLearned: A Foundation Model Framework for All Tasks Involving Jet Physics

### v1

@@ I Introduction
Hadronic jets are the result of high-energy quarks and gluons. These objects are ubiquitous in reactions at colliders and elsewhere that involve momentum transfers well above 1 GeV. The pattern of the radiation pattern within jets encodes information about their origin as well as emergent properties of the strong force. For decades, researchers proposed physically-inspired observables to characterize this complex structure [1, 2, 3]. Deep learning-based taggers are now standard, but it is challenging to assemble enough examples jets to train state-of-the-art machine learning models for each of the myriad tasks involving jets.
@@
In this paper, we present a significant upgrade of OmniLearn. This includes a number of updates to the architecture and training to improve performance. These updates also allow for more per-jet-constituent information to be incorporated into the model as well as for datasets without labels to be used in the foundation model training. The upgraded model, called OmniLearned, is now trained with over one billion jets (10x the previous model). This is a significant milestone on its own and utilizes the scaling properties of transformers. Alongside this paper, we have prepared software for readily accessing all of the train and test datasets in a unified framework as well as well-documented code for training/fine-tuning the machine learning models. The new framework is demonstrated on three key tasks. First, the models are compared with the current state of the art on the community benchmark top quark and quark-versus-gluon tagging datasets. Second, we are able to improve the performance of $b$-tagging with ATLAS public simulations. As part of this task, we also show how to repurpose parts of the foundation model for tasks different than their original intention (per-track classification from jet generator). Lastly, we show how the top quark cam be effectively re-discovered using anomaly detection deployed on CMS open data. Along the way, we show how the outputs of the foundation model can also be used directly (without fine tuning) for anomaly detection when the target is relatively close to a class in the training dataset.
@@ II.1 Input Features
In the previous OmniLearn implementation, the inputs to the model were the kinematic information of each particle clustered in the jet. While only four numbers are necessary to completely describe the momentum of each particle, redundant features were added to accelerate training convergence. Such features included the difference in pseudorapidity $\eta$ and azimuthal angle $\phi$ between each particle and the jet axis, together with the absolute distance $\Delta_{R}$, calculated using the same $\eta$ and $\phi$ values. In OmiLearned, we maintain the difference in $\eta$ and $\phi$ between the particle and jet axis as inputs but only add two additional features, consisting of the logarithm of the transverse momentum $\log p_{T}$ and energy $\log E$. We also remove the pre-processing step, previously used to standardize the mean and variance of the input features, since this choice of representation already leads to each input roughly centered at zero with $\mathcal{O}(1)$ variance.
@@ II.6 Loss Function
$=\textsc{CE}(y,y_{\text{pred}})+\left\|\mathbf{v}-\mathbf{v}_{\text{pred}}\right\|^{2}+\alpha^{2}\textsc{CE}(y,\hat{y}_{\text{pred}})\,.$ (3)
@@
We also modify the training of the generative part based on diffusion models. During the diffusion model training, particles clustered inside the jet were first perturbed using a time-dependent Gaussian value::
@@
Even though both formulations lead to similar losses, the training convergence rate and sample quality between diffusion and flow matching models can be different, with flow matching models often leading to better generation quality with similar number of sampling steps as the diffusion training. For this reason, we change the velocity prediction from OmniLearn to the flow matching objective in Eq. 6. This change also leads to the modification to the function $\alpha$ in Eq. 3 that multiplies the cross entropy term evaluated over perturbed inputs.
@@ IV Jet Classification
The last classification task we investigate uses the dataset released by the ATLAS Collaboration for flavor tagging [69]. Smaller radius jets with R = 0.4 are clustered using the anti-$k_{t}$ algorithm and simulated using top quark decays from POWHEG BOX [70, 71, 72, 73] interfaced with Pythia 8 at next-to-leading-order accuracy. Reconstructed tracks are used as inputs for the training. Jets are matched to hadron-level jets if they are found in a radius of 0.3 and are assigned the parton flavor as a label. Four different classes are considered for the classification task: b(ottom)-jets, c(harm)-jets, light jets, and jets produced through hadronic decays from tau leptons. Besides the basic kinematic quantities used by OmniLearned, additional vertex information is available. To promote a fair comparison with current flavor tagging algorithms used by the ATLAS Collaboration, we include all features that are also used by the Collaborations’s current state if the art GN2 architecture, with detailed explanation for each feature presented in Ref. [69]. In addition to the main classification task, GN2 also includes auxiliary tasks that help improve the performance of the tagger. These include the prediction of the track origin and the identification of tracks sharing a common vertex. The first auxiliary task assigns a label to each track, similar to the PID information present in other datasets. When training OmniLearned, we perform the jet classification using the classification but re-purpose the generation head to perform track classification, thus enabling the model to benefit from the additional auxiliary task. In this case, all the weights from the generation head are loaded during the fine-tuning process while the output layers are replaced to match the number of classes present in each classification task. Even though the diffusion task is different than the track classification we observe the use of the pre-trained weights to still be beneficial compared to starting from random weights.
@@
Table 4: Comparison between the performance reported for different classification algorithms on the ATLAS flavor tagging dataset. Values reported fo GN2 are obtained from the model trained on the same dataset for a fair comparison. While the results of a single model trained are displayed, we notice that the performance listed is stable against multiple runs of the same algorithms apart from GN2 whose baseline is taken from the public results.
@@ V Anomaly Detection
For all model sizes the signal component is clearly visible, with the medium size yielding again the highest sensitivity. Compared to the CATHODE strategy, the observed significance is higher. This application directly uses OmniLearned without additional fine-tuning, resulting in faster deployment.
@@ VI Conclusion and Outlook
The code for this paper can be found at https://github.com/ViniciusMikuni/OmniLearned.
@@ Acknowledgments
We thank Pradyun Hebbar, Joschka Birk, Runze Li, and Ibrahim Elsharkawy for helpful comments and testing the software environment during development. We thank our colleagues from the H1 Collaboration for allowing us to use the simulated MC event samples. We also thank DESY-IT and the MPI für Physik for providing computing infrastructure and supporting the data preservation project of the HERA experiments. VM is supported by the EXPERT-J program as part of the J-RISE Initiative No. EX25-009-01. BN is supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-76SF00515. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0021099 and HEP-ERCAP0028249.

### v2

# Foundation Model Framework for All Tasks Involving Jet Physics

@@ I Introduction
Hadronic jets are the result of high-energy quarks and gluons. These objects are ubiquitous in reactions at colliders and elsewhere that involve momentum transfers well above 1 GeV. The radiation pattern within jets encodes information about their origin as well as emergent properties of the strong force. For decades, researchers proposed physically-inspired observables to characterize this complex structure [1, 2, 3]. Deep learning-based taggers are now standard, but it is challenging to assemble enough examples jets to train state-of-the-art machine learning models for each of the myriad tasks involving jets.
@@
In this paper, we present a significant upgrade of OmniLearn. This includes a number of updates to the architecture and training to improve performance. These updates also allow for more per-jet-constituent information to be incorporated into the model as well as for datasets without labels to be used in the foundation model training. The upgraded model, called OmniLearned, is now trained with over one billion jets (10x the previous model). This is a significant milestone on its own and utilizes the scaling properties of transformers. Alongside this paper, we have prepared software for readily accessing all of the train and test datasets in a unified framework as well as well-documented code for training/fine-tuning the machine learning models. The new framework is demonstrated on three key tasks. First, the models are compared with the current state of the art on the community benchmark top quark and quark-versus-gluon tagging datasets. Second, we are able to improve the performance of $b$-tagging with ATLAS public simulations. As part of this task, we also show how to repurpose parts of the foundation model for tasks different than their original intention (per-track classification from jet generator). Lastly, we show how the top quark can be effectively re-discovered using anomaly detection deployed on CMS open data. Along the way, we show how the outputs of the foundation model can also be used directly (without fine tuning) for anomaly detection when the target is relatively close to a class in the training dataset.
@@ II.1 Input Features
In the previous OmniLearn implementation, the inputs to the model were the kinematic information of each particle clustered in the jet. While only four numbers are necessary to completely describe the momentum of each particle, redundant features were added to accelerate training convergence. Such features included the difference in pseudorapidity $\eta$ and azimuthal angle $\phi$ between each particle and the jet axis, together with the absolute distance $\Delta_{R}$, calculated using the same $\eta$ and $\phi$ values. In OmiLearned, we maintain the difference in $\eta$ and $\phi$ between the particle and jet axis as inputs but only add two additional features, consisting of the logarithm of the transverse momentum $\log p_{T}$ and energy $\log E$. We also remove the pre-processing step, previously used to standardize the mean and variance of the input features, since this choice of representation already leads to each input roughly centered at zero with $\mathcal{O}(1)$ variance. That leaves us with the minimal set of inputs supported by OmniLearned as $(\Delta\eta,\Delta\phi,\log p_{T},\log E)$.
@@ II.6 Loss Function
$=\textsc{CE}(y,y_{\text{pred}})+\left\|v-v_{\text{pred}}\right\|^{2}+\alpha^{2}\textsc{CE}(y,\hat{y}_{\text{pred}})\,.$ (3)
@@
We also modify the training of the generative part based on diffusion models. During the diffusion model training, particles clustered inside the jet were first perturbed using a time-dependent Gaussian value:
@@
Even though both formulations lead to similar losses, the training convergence rate and sample quality between diffusion and flow matching models can be different, with flow matching models often leading to better generation quality with similar number of sampling steps as the diffusion training. For this reason, we change the velocity prediction from OmniLearn to the flow matching objective in Eq. 6 resulting in the combined loss:
$\mathcal{L}=\textsc{CE}(y,y_{\text{pred}})+\left\|u-u_{\text{pred}}\right\|^{2}+\alpha^{2}\textsc{CE}(y,\hat{y}_{\text{pred}})\,.$ (7)
This change also leads to the modification to the function $\alpha$ in Eq. 3 that multiplies the cross entropy term evaluated over perturbed inputs.
@@ IV Jet Classification
The last classification task we investigate uses the dataset released by the ATLAS Collaboration for flavor tagging [69, 70]. Smaller radius jets with R = 0.4 are clustered using the anti-$k_{t}$ algorithm and simulated using top quark decays from POWHEG BOX [71, 72, 73, 74] interfaced with Pythia 8 at next-to-leading-order accuracy. Reconstructed tracks are used as inputs for the training. Jets are matched to hadron-level jets if they are found in a radius of 0.3 and are assigned the parton flavor as a label. Four different classes are considered for the classification task: b(ottom)-jets, c(harm)-jets, light jets, and jets produced through hadronic decays from tau leptons. Besides the basic kinematic quantities used by OmniLearned, additional vertex information is available. To promote a fair comparison with current flavor tagging algorithms used by the ATLAS Collaboration, we include all features that are also used by the Collaboration’s current state if the art GN2 architecture, with detailed explanation for each feature presented in Ref. [69]. In addition to the main classification task, GN2 also includes auxiliary tasks that help improve the performance of the tagger. These include the prediction of the track origin and the identification of tracks sharing a common vertex. The first auxiliary task assigns a label to each track, similar to the PID information present in other datasets. When training OmniLearned, we perform the jet classification using the classification but re-purpose the generation head to perform track classification, thus enabling the model to benefit from the additional auxiliary task. In this case, all the weights from the generation head are loaded during the fine-tuning process while the output layers are replaced to match the number of classes present in each classification task. Even though the diffusion task is different than the track classification we observe the use of the pre-trained weights to still be beneficial compared to starting from random weights.
@@
Table 4: Comparison between the performance reported for different classification algorithms on the ATLAS flavor tagging dataset. Values reported for GN2 are obtained from the model trained on the same dataset for a fair comparison. While the results of a single model trained are displayed, we notice that the performance listed is stable against multiple runs of the same algorithms apart from GN2 whose baseline is taken from the public results.
@@ V Anomaly Detection
For all model sizes the signal component is clearly visible. However, bigger model sizes lead to a small degradation of the model sensitivity. This is likely due to imperfect simulation details used during the pre-training. This application directly uses OmniLearned without additional fine-tuning, resulting in faster deployment.
@@ VI Conclusion and Outlook
The code and data for this paper can be found at [99] https://github.com/ViniciusMikuni/OmniLearned.
@@ Acknowledgments
We thank Pradyun Hebbar, Joschka Birk, Runze Li, and Ibrahim Elsharkawy for helpful comments and testing the software environment during development. We also thank Kevin Grief for discussion on the use of the ATLAS flavor tagging dataset. We thank our colleagues from the H1 Collaboration for allowing us to use the simulated MC event samples. We also thank DESY-IT and the MPI für Physik for providing computing infrastructure and supporting the data preservation project of the HERA experiments. VM is supported by JST EXPERT-J, Japan Grant Number JPMJEX2509. BN is supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-76SF00515. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards ERCAP0034229, HEP-ERCAP0021099 and HEP-ERCAP0028249.